OpenStudy (anonymous):

f'(x)=(2 ln(x))/x How do I find the x value? I put the equation equal to zero, but I do not know any log rule to solve it for x value. I need help please.

OpenStudy (anonymous):

Do you want to solve the equation for x?

OpenStudy (anonymous):

Yes

OpenStudy (across):

That's not possible, unless you're setting f'(x)=0.

OpenStudy (across):

Are you trying to find maxima/minima?

OpenStudy (across):

\[\frac{2\ln(x)}{x}=0\]doesn't seem too bad.

OpenStudy (anonymous):

if you are looking for critical value , look for where function diverge too

OpenStudy (anonymous):

How do I proceed after this to find the value for x?

OpenStudy (anonymous):

Yes I am looking for critical values

OpenStudy (anonymous):

what happened if you plug in 0 for x

OpenStudy (anonymous):

you can't take the ln(0). There has to be some other way related to log rules to solve this question, and I can't figure it out.

OpenStudy (anonymous):

it diverges, which mean it is a criticle point. a critical point of a function of a real variable is any value in the domain where either the function is not differentiable or its derivative is 0

OpenStudy (across):

\[\frac{2\ln(x)}{x}=0,\]\[2\ln(x)=0,\]\[\ln(x)=0,\]\[x=1.\]